# Fine-Grained Domain Generalization with Feature Structuralization

### v2

@@ IV-B Implementation and Evaluation
The feature extractor $E$ can be instantiated with ResNet (RN) series [51], Vision Transformer(ViT) series [52], and ASMLP series [53] backbones, respectively. The learning rate is initially set to 0.003 and follows a dynamic strategy, where the coefficient of the learning rate varies from 1 to 0.1 during the training process. All layers except the backbones are trained from scratch, and their learning rate is 10 times that of the backbone. The model is optimized using mini-batch SGD with a momentum of 0.9. A single 3090 GPU with a batch size fixed to 32 is used for training and evaluation. Typically, the coefficients are configured as $\lambda_{cs}=0.05$, $\lambda_{cd}=0.5$, and $\lambda_{p}=1$. $r^{c}$ and $r^{p}$ are set to 0.5 and 0.3, respectively. We employ fine-grained classification accuracy as the evaluation criterion. All models are trained three times and evaluated for ten trials on the target domain each time. Due to FSDG retaining only the finest branch during inference, the number of parameters of the dual backbone model is the same as that of the single backbone model, which is 24M.
@@ IV-C Main Results
(a) Similarity Value $S_{cs}$
(b) Similarity Value $S_{cd}$
(c) Similarity Value $S_{p}$
@@ IV-D Analysis
Ablation study on the losses. We present the outcomes of models optimized under various combinations of losses. As shown in Table V, each loss and their combinations contribute to performance improvement compared to the baseline FGDG model (the first row). The efficacy of the FS approach is substantiated by quantitative assessments of FGDG performance.
@@
Ratio impact. FSDG reaches its peak when the proportions of commonality, specificity, and confounding are controlled to be 5:3:2, as Fig. 4 shows. As the ratios of commonality and specificity decrease, the performance in FGDG drops to a minimum. Compared to specificity, commonality has a greater impact on generalization. The constraints on the common and specific parts contribute to the improved generalization. Even the lowest performance (54.48%) at the point where common and specific components only account for 20% each is still higher than the case without the $\mathcal{L}_{FS}$ constraint, which is 53.66% in Table I. This is because, in this study, we impose no strict constraints on the confounding part, which means increasing the ratio of the confounding part tends to resemble the original FGDG model. Moreover, performance is degraded when $r^{n}$ is configured as 0, highlighting the necessity of isolating the confounding part from the commonality and specificity at the feature level.
@@
TABLE VIII: 4-granularity labels of 8 category examples and the overlap of statistical top concepts between each category and the other 7 categories. The ratio of commonality shows the ratio between the Common and the All.
@@ V Explainability Analysis
In this paper, we enhance the FGDG performance through the FS method with the integration of data and knowledge. The demonstrated enhancement primarily manifests in that the common part of the features, decoupled based on channel indices, primarily focuses on the commonalities and representing the invariance between categories, while the specific part focuses on learning discriminative characteristics. In this section, we endeavor to validate that FS increases the explicit concept matching intensity between the shared concepts among categories and the indices of model channels.
We argue that categories that are closer together should activate more identical concepts in the common part, while classes that are farther apart activate fewer shared concepts in this segment. To demonstrate this argument, we randomly select a sequence of classes from the Cub-Paintings dataset, and their corresponding category labels at four granularities are shown in Table VIII. We then define distances based on a multi-granular structure to roughly measure the similarity between fine-grained categories:
@@
$\mathcal{S}_{class}$ is a discrete numerical value, with its maximum determined by the number of granularities in the hierarchical knowledge structure. $\mathcal{S}_{class}$ provides a ranking of the differences between different fine-grained categories. It should be noted that $\mathcal{S}_{class}=0$ does not imply that there is no commonality between two categories but rather indicates low similarity between them within the multi-granular structure. For example, in the Cub-Paintings dataset, all bird species share a consistent body shape, which is a form of commonality. However, this similarity might not be reflected within the current granularity structure. Fig. 5a displays the confusion matrix of $\mathcal{S}_{class}$ between fine-grained category instances as listed in Table VIII. Fig. 5a depicts the ground truth quantifying the difference ranks between categories.
@@
Fig. 5b-5d illustrate the distributions of the top 26 relevant concepts in three parts. We find that FS leads to an increase in the number of concepts activated in the common region while reducing the activation in the other two regions. For example, among the top 26 concepts in category 10, 22 over 26 concepts are activated at the common part by FS, whereas the baseline FGDG model only has 11. This suggests that, for fine-grained recognition, extracting common concepts is more crucial for generalization as common concepts reflect the invariance between categories.
Intuitively, we argue that closer categories should share more common concepts, meaning that there should be a higher overlap in the most relevant activated channels. We compute the confusion matrix of the overlap of relevant concepts between each pair of category examples, as illustrated in Fig. 6. Compared among Fig. 5a, 6a and 6e, the overall shape of the concept overlap activated by FS closely resembles the ground truth. The FGDG model, without FS, overlooks the relationships between categories, resulting in distant classes having a considerable overlap in activated shared concepts, which is inaccurate. Besides, as shown in Fig. 6b and 6f, FS leads to greater activation of shared concepts in the common part. Besides, the shared concepts between categories are more concentrated in the common partition, which aligns with our initial assumptions and objectives. This also indicates that we, to some extent, pre-define the functionalities of the channels, enhancing the model’s internal transparency and explainability.
@@
When contrasting the distribution of shared concepts in the specific region, as shown in Fig. 6d and Fig. 6h, we observe that FS elicits a sparser distribution of shared concepts. This indicates that this region focuses more on specificities, aligning with the initial assumptions of the model. This phenomenon also occurs in the confounding part. Since we do not impose constraints on the confounding part, a natural discussion arises as to whether applying constraints only to the common and specific parts would increase more activations in the confounding part. In contrast, by analyzing Fig. 5d, Fig. 6d and Fig. 6h, we find that the number of top activations in the confounding part is decreased. All of these results demonstrate that FS directs the model to focus more on capturing the relationships between categories rather than on other spurious correlations, therefore improving generalization.
To quantitatively analyze the matching intensity between shared concepts and channel structure regions, we sum up the number of shared concepts between each category and the other seven categories, as shown in Table VIII. By calculating the proportion of shared concepts within the common region relative to all shared concepts, we find that FS increases the explicit matching intensity between shared concepts and model channels from 40% to 68%. Therefore, our FS approach is indeed effective.

### v3

@@ IV-B Implementation and Evaluation
The feature extractor $E$ can be instantiated with ResNet (RN) series [57], Vision Transformer(ViT) series [58], and ASMLP series [59] backbones, respectively. The learning rate is initially set to 0.003 and follows a dynamic strategy, where the coefficient of the learning rate varies from 1 to 0.1 during the training process. All layers except the backbones are trained from scratch, and their learning rate is 10 times that of the backbone. The model is optimized using mini-batch SGD with a momentum of 0.9. A single 3090 GPU with a batch size fixed to 32 is used for training and evaluation. A progressive grid search strategy is employed to identify the optimal weights for our proposed optimization objectives. The search space is configured as a set [0.005, 0.01, 0.05, 0.1, 0.5, 1]. Typically, the coefficients are configured as $\lambda_{cs}=0.05$, $\lambda_{cd}=0.5$, and $\lambda_{p}=1$. $r^{c}$ and $r^{p}$ are set to 0.5 and 0.3, respectively. We employ fine-grained classification accuracy as the evaluation criterion. All models are trained three times and evaluated for ten trials on the target domain each time. Due to FSDG retaining only the finest branch during inference, the number of parameters of the dual backbone model is the same as that of the single backbone model, which is 24M.
@@ IV-D Analysis
Ablation study on the losses. We present the outcomes of models optimized under various combinations of losses. As shown in Table V, each loss and their combinations contribute to performance improvement compared to the baseline FGDG and FGDG (+$\mathcal{L}_{lf}$) model (the first and second rows). Rows 3, 4, and 5 respectively demonstrate the increase in performance when each of the constraints is individually incorporated. Compared to row 2, which lacks feature partitioning and $\mathcal{L}_{dec}$, their inclusion results in generalization improvements of 0.85%, 0.52%, and 0.86%. When all three constraints act in concert (second-to-last row), the generalization is enhanced by 1.44%. The contrast between row 6 and row 2 reveals that the sole introduction of $\mathcal{L}_{dec}$ yields a 0.43% improvement in generalization. When $\mathcal{L}_{dec}$ is optimized in synergy with the commonality and specificity constraints (last row), the model’s generalization is elevated to 55.65%, marking a 1.99% improvement. The addition of $\mathcal{L}_{dec}$ to the combined commonality and specificity constraints (second-to-last row) further boosts generalization by an additional 0.55%. These results all demonstrate the importance and effectiveness of the FS method proposed in this paper.
(a) Similarity Value $S_{cs}$
(b) Similarity Value $S_{cd}$
(c) Similarity Value $S_{p}$
@@
Ratio impact. FSDG reaches its peak when the proportions of commonality, specificity, and confounding are controlled to be 5:3:2, as Fig. 4 shows. As the ratios of commonality and specificity decrease, the performance in FGDG drops to a minimum. Compared to specificity, commonality has a greater impact on generalization. The constraints on the common and specific parts contribute to the improved generalization. Even the lowest performance (54.48%) at the point where common and specific components only account for 20% each is still higher than the case without the $\mathcal{L}_{FS}$ constraint, which is 53.66% in Table I. This is because, in this study, we impose no strict constraints on the confounding part, which means increasing the ratio of the confounding part tends to resemble the original FGDG model. Moreover, performance is degraded when $r^{n}$ is configured as 0, highlighting the necessity of isolating the confounding part from the commonality and specificity at the feature level. These results all validate the necessity of structuring features into the three segments.
@@
TABLE VIII: Classification accuracy (%) of FGVC methods on FGDG task.
      Method       C→P       P→C       Avg
      MPSA [61]       59.07       36.31       47.69
      SEF [62]       58.09       35.63       46.86
TABLE IX: Classification accuracy (%) of in-domain and out-of-domain scenarios.
Method In-Domain Acc. Out-of-Domain Acc. (C→P)
S-FGDG (Baseline) 87.71 61.21
S-FGDG (+$\mathcal{L}_{lf}$) 86.98 61.98
S-FSDG (Ours) 86.37 63.42
@@
Comparison with FGVC methods. We conduct experiments to evaluate FGVC methods on the domain generalization task, with the results presented in Table VIII. The generalization capability of FGVC methods is consistently lower than that of the proposed FSDG model. Compared to the generalization of the single-backbone S-FSDG (54.14%), the generalization of the FGVC method is 6.87% lower. In pursuing higher fine-grained recognition accuracy, they inadvertently learn spurious correlations, which compromise their generalization performance. When sample distributions shift, such as changes in background or style, the recognition accuracy of these methods decreases significantly.
(a) Single optimization
(b) Progressive optimization
Figure 5: Performance of various loss functions under different coefficients. (a) presents the performance of each loss. (b) is the progressive optimization of the 4 losses.
Evaluation on in-domain and out-of-domain scenarios. We conduct the evaluation of the FSDG and baseline models on both in-domain and out-of-domain settings, as detailed in Table IX. Three models are trained on the CUB-200-2011 dataset, with in-domain performance assessed on the validation subset and out-of-domain performance on the CUB-200-Painting domain. The results indicate that the FSDG model experiences a performance decrement of approximately 1.34% in the in-domain setting while demonstrating a significant enhancement of 2.21% in the out-of-domain setting. This outcome underscores that the enhancement in generalization brought about by FS is directly attributed to the model’s inherent generalization capabilities. It effectively mitigates the influence of spurious correlations such as background and style, enabling the model to learn domain-invariant features.
Analysis of coefficients on loss functions. We first conduct an optimal coefficient search for each loss. As shown in Fig. 5a, each loss possesses an optimal value within the search space and confers an enhancement in generalization. Starting at $\lambda_{cs}$, we then progressively identify the optimal coefficient and fix it to search for the optimal point of the subsequent losses. As depicted in Fig. 5b, this strategy results in a steady improvement of the model’s performance, culminating in the optimal generalization performance (55.65%). Ultimately, the optimal set of parameters is obtained.
TABLE X: 4-granularity labels of 8 category examples and the overlap of statistical top concepts between each category and the other 7 categories. The ratio of commonality shows the ratio between the Common and the All.
@@ V Explainability Analysis
In this paper, we enhance the FGDG performance through the FS method with the integration of data and knowledge. The demonstrated enhancement primarily manifests in that the common part of the features, decoupled based on channel indices, primarily focuses on the commonalities and representing the invariance between categories, while the specific part focuses on learning discriminative characteristics. In this section, we endeavor to validate that FS increases the explicit concept matching intensity between the shared concepts among categories and the indices of model channels, and the FSDG model embeds knowledge of a multi-granularity structure within its feature space.
We argue that categories that are closer together should activate more identical concepts in the common part, while classes that are farther apart activate fewer shared concepts in this segment. To demonstrate this argument, we randomly select a sequence of classes from the Cub-Paintings dataset, and their corresponding category labels at four granularities are shown in Table X. We then define distances based on a multi-granular structure to roughly measure the similarity between fine-grained categories:
@@
$\mathcal{S}_{class}$ is a discrete numerical value, with its maximum determined by the number of granularities in the hierarchical knowledge structure. $\mathcal{S}_{class}$ provides a ranking of the differences between different fine-grained categories. It should be noted that $\mathcal{S}_{class}=0$ does not imply that there is no commonality between two categories but rather indicates low similarity between them within the multi-granular structure. For example, in the Cub-Paintings dataset, all bird species share a consistent body shape, which is a form of commonality. However, this similarity might not be reflected within the current granularity structure. Fig. 6a displays the confusion matrix of $\mathcal{S}_{class}$ between fine-grained category instances as listed in Table X. Fig. 6a depicts the ground truth quantifying the difference ranks between categories.
@@
To quantitatively analyze the matching intensity between shared concepts and channel structure regions, we sum up the number of shared concepts between each category and the other seven categories, as shown in Table X. By calculating the proportion of shared concepts within the common region relative to all shared concepts, it can be observed that the ratio of overlapping concepts that fall within the commonality segment for models without FS is 40% (even lower than the 50% ‘common’ ratio of the partitioning scheme based on a 5:3:2 ratio). Upon the incorporation of the FS method, this proportion increases to 68%. The proportion of these overlapping concepts within the specificity segment decreases (from 36% to 17%), indicating an increased focus on category-specific concepts in this segment. Besides, Fig. 6b-6d illustrate the distributions of the top 26 relevant concepts in three parts. FS leads to an increase in the number of concepts activated in the common region while reducing the activation in the other two regions. These findings validate that FS successfully confines the semantical commonalities among categories to the first segment, while the second segment recognizes concepts unique to each category. When predicting the given samples, the FSDG model pays more attention to the general semantics through its common segment.
TABLE XI: Spearman Rank Correlation Coefficient of concept overlap between regional features of various models and the ground truth. The second column represents the value between all features. The third column is the comparison between ground truth and the common segment.
FGDG All Com. Spe. Conf.
Ground Truth 0.62 0.52 0.55 0.67
FSDG All Com. Spe. Conf.
Ground Truth 0.88 0.83 0.50 0.60
We further substantiate that the FSDG model embeds a multi-granularity structure into its feature embeddings, enabling it to perceive the similarities between categories. Intuitively, we argue that closer categories should share more common concepts, meaning that there should be a higher overlap in the most relevant activated concepts. We compute the confusion matrix of the overlap of relevant concepts between each pair of category examples, as illustrated in Fig. 7. Considering that the FGDG model does not undergo partitioned training, we compare the learning capabilities of the FGDG and FSDG models with respect to the multi-granularity structure across all features (i.e. Fig. 7a and Fig. 7e). Compared to Fig. 7a, the overall shape of the concept overlap activated by FS (Fig. 7e) closely resembles the ground truth (Fig. 6a). We proceed to calculate the Spearman Rank Correlation Coefficient for the overlapping concepts between different segments and the ground truth for both models, as presented in Table XI. It is observed that the FSDG model achieves a correlation coefficient of 0.88 in learning explicit category similarity, whereas the FGDG model only reaches 0.62. The common segment of the FSDG model also exhibits a correlation coefficient of 0.83, enabling the learning of general features between categories. The correlation between the other two segments and the ground truth is relatively weak. The FGDG model, without FS, overlooks the relationships between categories, resulting in distant classes having a considerable overlap in activated shared concepts, which is inaccurate. These results demonstrate that the FSDG model embeds a perception of multi-granularity structure within the feature space. When identifying objects, it is capable of perceiving the distances between fine-grained categories.
@@
When contrasting the distribution of shared concepts in the specific region, as shown in Fig. 7g, we observe that FS elicits a sparser distribution of shared concepts. The correlation between these features and the multi-granularity category similarity ground truth is only 0.5. This indicates that this region focuses more on specificities, aligning with the initial assumptions of the model. This phenomenon also occurs in the confounding part. Besides, we do not impose constraints on the confounding part, a natural discussion arises as to whether applying constraints only to the common and specific parts would increase more activations in the confounding part. In contrast, by analyzing Fig. 6d and Fig. 7h, we find that the number of top activations in the confounding part is decreased. All of these results demonstrate that FS embeds a multi-granularity structure into the feature space and achieves feature functionalization in accordance with the objectives. It directs the model to focus more on capturing general and specific features between categories, therefore improving generalization. This also indicates that we, to some extent, pre-define the functionalities of the channels, enhancing the model’s internal transparency and explainability.